A convention manger finds that she has $1540, made up of twenties and fifties. She has a total of 50 bills. How may fifty-dollar bills does the manager have?
Answer by JulietG(1812)

If all 50 bills were $50s, her total would be 50*50=2500
However, her total is 1540. 2500-1540=960
The difference between fifties and twenties is $30.
960 / 30 = 32
32 bills are $20. 32 * 20 = $640
If 32 are twenties, then the remaining 18 are fifties.
18 * $50 = $900
$900 + $640 = $1540